OpenStudy (anonymous):

Evalute 2n2 − 7n − 4 given n(n − 2) = 15.

OpenStudy (anonymous):

n^2 - 2n - 15 = 0 (n-5)(n+3) = 0 n = 5 or - 3 2n2 − 7n − 4 = 50 - 35 -4 = 11 (when n = 5) or 2n2 − 7n − 4 = 18 + 21 - 4 = 35

OpenStudy (anonymous):

I need a larger number and smaller number